# Device-Native Autonomous Agents for Privacy-Preserving Negotiations

### v1

@@ Abstract
Automated negotiations in insurance and business-to-business (B2B) commerce encounter substantial challenges. Current systems force a trade-off between convenience and privacy by routing sensitive financial data through centralized servers, increasing security risks, and diminishing user trust. This study introduces a device-native autonomous Artificial Intelligence (AI) agent system for privacy-preserving negotiations. The proposed system operates exclusively on user hardware, enabling real-time bargaining while maintaining sensitive constraints locally. It integrates zero-knowledge proofs to ensure privacy and employs distilled world models to support advanced on-device reasoning. The architecture incorporates six technical components within an agentic AI workflow. Agents autonomously plan negotiation strategies, conduct secure multi-party bargaining, and generate cryptographic audit trails without exposing user data to external servers. The system is evaluated in insurance and B2B procurement scenarios across diverse device configurations. Results show an average success rate of 87%, a 2.4 latency improvement over cloud baselines, and strong privacy preservation through zero-knowledge proofs. User studies show 27% higher trust scores when decision trails are available. These findings establish a foundation for trustworthy autonomous agents in privacy-sensitive financial domains.
@@ I Introduction
Autonomous agents present a promising direction to address these limitations [yao2023react, autogpt2023]. In bilateral negotiations, buyer and seller agents conduct real-time bargaining on behalf of their principals. Buyers can specify constraints, such as a maximum budget and acceptable trade-offs, while the agents manage the negotiation process automatically. However, several significant challenges persist: privacy (users must not disclose financial constraints to external servers), explainability (regulatory requirements mandate decision transparency [goodman2017european]), resource limitations (restricting on-device algorithm complexity), and fairness (preventing exploitative agreements [nash1950bargaining]).
This work introduces a device-native agentic AI system for autonomous negotiations. The proposed approach converts static pricing into privacy-preserving bargaining through a six-layer architecture that integrates the following technical contributions.
(1) Selective state transfer ensures continuity across devices.
(2) Explainable memory generates cryptographic audit trails.
(3) World model distillation facilitates on-device reasoning [hinton2015distilling].
(4) A privacy-preserving protocol secures multi-agent bargaining through the use of zero-knowledge proofs [groth2016size].
(5) Model-aware offloading optimizes the balance between latency and privacy.
(6) Simulation-critic safety mechanisms are implemented to prevent unfair outcomes.
@@ II-A Agentic AI Frameworks
Recent frameworks have advanced autonomous AI reasoning. ReAct [yao2023react] integrates reasoning and acting within language models. AutoGPT [autogpt2023] sequences large language model (LLM) calls to address complex tasks. LangChain [langchain2023] provides tools for constructing agent applications. Although these systems demonstrate autonomous capabilities, they share limitations. All depend on cloud infrastructure for execution and lack privacy-preserving mechanisms. None supports bilateral negotiation protocols or runs on resource-constrained devices.
The proposed system adopts a fundamentally different design philosophy by prioritizing privacy-first, on-device execution. Negotiation is treated as a core capability rather than an optional extension. The architecture operates on consumer hardware while preserving agent autonomy.
@@ II-B Multi-Agent Systems and Negotiation
Game theory provides foundational principles for automated negotiation. Nash bargaining [nash1950bargaining] defines fair outcomes in bilateral settings. Rubinstein’s alternating offers model [rubinstein1982perfect] describes sequential bargaining dynamics. Auction mechanisms [krishna2009auction] handle multi-party price discovery. These theoretical frameworks inform the protocol design presented in this work.
Recent multi-agent systems demonstrate potential for agent cooperation. MetaGPT [hong2023metagpt] assigns distinct roles to collaborating agents, while CAMEL [li2023camel] facilitates communication through role-playing mechanisms. However, these systems primarily address collaborative task completion rather than bilateral negotiation involving competing interests. Privacy considerations are minimally addressed in their designs.
Zero-knowledge proofs facilitate private computation without disclosing underlying data [goldwasser1989knowledge]. zk-SNARKs [groth2016size] enable efficient proof generation suitable for practical deployment. Recent research applies these cryptographic techniques to machine learning [kang2022scaling]. The present work integrates zero-knowledge proofs into negotiation protocols, enabling agents to demonstrate constraint satisfaction without revealing the constraints themselves.
@@ II-C Explainability and Trust
Research in explainable AI demonstrates that audit trails improve user trust [ribeiro2016should]. LIME and SHAP [lundberg2017unified] provide post-hoc explanations for model predictions. Financial AI systems must comply with regulations that require decision transparency [goodman2017european].
@@
Table I summarizes the comparison. The proposed framework combines autonomous reasoning, privacy-preserving negotiation, and cryptographic explainability within a device-native architecture. This combination addresses gaps unaddressed by existing systems.
System Capabilities Auto. Priv. Neg. Expl. Attest. Device State AutoGPT [autogpt2023] ✓ – – – – – – LangChain [langchain2023] ✓ – – – – – MetaGPT [hong2023metagpt] ✓ – – – – – CAMEL [li2023camel] ✓ – – – – – – Proposed Model ✓ ✓ ✓ ✓ ✓ ✓ ✓
@@ III-C Component 2: Explainable Memory
The explainable memory system produces cryptographic audit trails for each decision. Each record contains timestamp, decision type, inputs, reasoning trace, and outcome. These records are linked using Merkle trees [merkle1987digital], and the root hash is periodically anchored to a public blockchain to enable independent verification.
@@ III-D Component 3: World Model Distillation
Knowledge distillation with soft targets is applied [hinton2015distilling]:
@@ III-E2 Negotiation Execution
Zero-knowledge proofs are constructed using Groth16 zk-SNARKs [groth2016size], which support succinct proofs with fast verification. The underlying circuit encodes constraint satisfaction, ensuring that the offered price meets , without revealing or . Proof generation requires 80 ms on high-end devices.
@@
Termination adheres to Nash bargaining principles [nash1950bargaining]. Offers are considered converged when , after which agents compute . This approach maintains privacy while achieving outcomes that closely approximate the Nash solution. Performance results are presented in Table V.
@@ III-H Integration and Data Flow
The architecture achieves its design objectives through component integration. Privacy is preserved by retaining sensitive data on-device. Explainability is supported by cryptographic memory producing tamper-evident audit trails. Efficiency results from selective offloading and model distillation. Safety is maintained through simulation-critic validation, which prevents harmful actions prior to execution.
@@ IV-A Experimental Setup
Datasets. 2 public datasets are used for evaluation. The Medical Cost Personal dataset [choi2018insurance] contains 1,338 records with premium values ranging from $1,121 to $63,770, including features such as age, BMI, smoking status, and region. The Supply Chain Shipment Pricing dataset [watsky2019supply] comprises 10,324 B2B transactions with unit prices and features including product category, shipment mode, and freight costs.
@@ IV-B Component Performance
State Transfer. Table II presents selective state transfer results across device tiers. High-end devices achieve 70% compression with a migration time of 80ms. Mid-range devices reach 78% compression at 120ms, while low-end devices compress to 85% but require 180ms. Task continuity exceeds 85% across all tiers, indicating that users can resume negotiations without substantial context loss.
@@
Explainable Memory. Table III presents the memory system overhead and its impact on user trust. Write latency averages 5.2ms per decision record. Read latency is 2.1ms for point queries. Proof generation for audit verification requires 12.3ms, and storage overhead is 340 bytes per decision.
@@
World Model Distillation. Table IV compares the performance of teacher and student models. The 7B-parameter teacher model achieves 94% accuracy in predicting negotiation outcomes with a cloud latency of 1000ms. The 500M-parameter student model maintains 88% accuracy with an on-device latency of 180ms for high-end devices.
@@
Model size reduction enables device deployment. The 6% accuracy loss is acceptable given 5.5 latency improvement on high-end devices.
Multi-Agent Negotiation. Table V presents the performance of the negotiation protocol. Simple scenarios achieve a 90% success rate with a fairness score of 0.88. Medium scenarios attain 87% success and a fairness score of 0.85. Complex scenarios demonstrate 85% success and a fairness score of 0.83.
@@ IV-C End-to-End Performance
Insurance Negotiation. Table VI summarizes results for high-end devices across three complexity levels. The system achieves 83-88% success rates with latency ranging from 420ms-740ms. Privacy leakage remains minimal (12-16 bits) compared to cloud approaches (256 bits).
@@
B2B Negotiation. Table VII presents B2B procurement results, showing 85-89% success rates with 380ms-680msms latency. B2B scenarios demonstrate marginally better performance than insurance, likely due to more structured negotiation parameters.
@@
Figure 3 shows success rate decline with complexity. Insurance exhibits steeper degradation (5 points) than B2B (4 points), reflecting its more contentious nature. Both domains maintain ¿83% success.
Figure 4 shows latency breakdown by component. The negotiation protocol accounts for 45-50% of total time, planning contributes 20-25%, and state management and safety checks comprise the remainder.
@@ IV-D Baseline Comparison
The proposed system achieves 2.4 speedup over Cloud-Only (420ms vs 1000ms) while Device-Only is fastest (350ms) but sacrifices success. Privacy is superior (14 bits leaked vs 256 for Cloud-Only and 180 for Naive Edge). Device-Only has zero leakage but cannot handle complex reasoning. Energy consumption is balanced at 5.0J (vs 2.0J for Cloud-Only, 12.0J for Device-Only).
@@ IV-E Ablation Study
State Transfer Removal. Latency increases 108% (875ms vs. 420ms) as full state must be transmitted. Success drops 3% due to timeout failures.
World Model Removal. Cloud-based calls replace on-device planning, resulting in a 119% increase in latency (920ms) and a 5% decrease in success rate. Privacy leakage also increases because planning data is transmitted to the cloud.
Negotiation Protocol Removal. Success rate decreases by 15% (72% vs 87%) as agents cannot verify constraint satisfaction or engage in multi-round convergence. Latency improves by 17% (349ms vs 420ms) because ZK proof generation and verification are eliminated, but this represents a loss of essential privacy-preserving functionality rather than a viable optimization.”
@@
Figure 6 illustrates the contributions of each component. State transfer enables mobility, the world model supports on-device reasoning, the negotiation protocol facilitates bargaining, memory establishes trust, offloading improves efficiency, and the safety critic prevents unfair outcomes.
@@ V Discussion
Key Insights. Device-native agents demonstrate effective negotiation performance with 87–90% success rates, matching or exceeding cloud-based alternatives while preserving privacy. Explainability increases user trust by 27% with minimal overhead (5ms per decision), consistent with prior research showing that transparency is crucial for AI systems in finance [wilson2025explainable]. This suggests that users highly value transparency when delegating financial decisions to autonomous agents.
The proposed architecture is applicable beyond insurance and B2B domains. Core technical requirements such as bilateral bargaining [nash1950bargaining, rubinstein1982perfect], privacy-sensitive constraints [goldwasser1989knowledge], and audit trails [goodman2017european] are also present in real estate negotiations, healthcare cost discussions, and legal settlements. The cryptographic protocol and on-device execution model are domain-agnostic, requiring only the definition of domain-specific constraints and the training of the world model [hinton2015distilling].
@@ VI Limitations
Experimental Scope. Our evaluation was conducted in controlled scenarios [choi2018insurance, watsky2019supply]. Real-world deployment studies are necessary to achieve more robust validation [ratliff2024framework].
System Constraints. The system currently supports only bilateral negotiations [nash1950bargaining, rubinstein1982perfect], and multi-party scenarios involving coalitions [back2024coalition] remain an area for future research. The protocol assumes cooperative bargaining, requiring additional safeguards to address adversarial counterparties. Device limitations also restrict the reasoning depth of the 500M-parameter model [hinton2015distilling] compared to cloud-scale models.
Behavioral Privacy Leakage. While the system achieves a 94% reduction in cryptographic data exposure, zero-knowledge proofs [goldwasser1989knowledge] do not prevent behavioral inference attacks [jin2024timing]. Adversaries observing negotiation dynamics may infer constraint proximity. Current metrics quantify only cryptographic leakage (Sec. IV). Mitigation strategies, such as differential privacy for timing [ratliff2024framework] and randomized concession schedules, are identified as directions for future work.
Offline Collusion. The system cannot prevent offline collusion between parties who coordinate externally and use agents to generate audit trails. This is a social rather than technical limitation. However, cryptographic audit trails [merkle1987digital] can facilitate post-hoc fraud detection [chassang2022detecting, tas2024machine] to support regulatory compliance [goodman2017european].

### v2

@@ Abstract
Automated negotiations in insurance and business-to-business (B2B) commerce encounter substantial challenges. Current systems force a trade-off between convenience and privacy by routing sensitive financial data through centralized servers, increasing security risks, and diminishing user trust. This study introduces a device-native autonomous Agentic AI system for privacy-preserving negotiations. The proposed system operates exclusively on user hardware, enabling real-time bargaining while maintaining sensitive constraints locally. It integrates zero-knowledge proofs to ensure privacy and employs distilled world models to support advanced on-device reasoning. The architecture incorporates six technical components within an Agentic AI workflow. Agents autonomously plan negotiation strategies, conduct secure multi-party bargaining, and generate cryptographic audit trails without exposing user data to external servers. The system is evaluated in insurance and B2B procurement scenarios across diverse device configurations. Results show an average success rate of 87%, a 2.4 reduction in latency relative to cloud baselines, and strong privacy preservation through zero-knowledge proofs. User studies show 27% higher trust scores when decision trails are available. These findings establish a foundation for trustworthy autonomous agents in privacy-sensitive financial domains.
@@ I Introduction
Autonomous agents present a promising direction to address these limitations [20, 15]. In bilateral negotiations, buyer and seller agents conduct real-time bargaining on behalf of their principals. Buyers can specify constraints, such as a maximum budget and acceptable trade-offs, while the agents manage the negotiation process automatically. However, several significant challenges persist: privacy (users must not disclose financial constraints to external servers), explainability (regulatory requirements mandate decision transparency [4]), resource limitations (restricting on-device algorithm complexity), and fairness (preventing exploitative agreements [13]).
This work introduces a device-native Agentic AI system for autonomous negotiations. The proposed approach converts static pricing into privacy-preserving bargaining through a six-layer architecture that addresses these challenges through (1) Selective state transfer, (2) Explainable memory, (3) World model distillation [6], (4) Privacy-preserving protocols with zero-knowledge proofs [5], (5) Model-aware offloading, and (6) Simulation-critic safety mechanisms.
@@ II-A Agentic AI Frameworks
Recent frameworks have advanced autonomous AI reasoning. ReAct [20] integrates reasoning and acting within language models. AutoGPT [15] sequences large language model (LLM) calls to address complex tasks. LangChain [1] provides tools for constructing agent applications. Although these systems demonstrate autonomous capabilities, they share limitations. All depend on cloud infrastructure for execution and lack privacy-preserving mechanisms. None supports bilateral negotiation protocols or runs on resource-constrained devices.
@@ II-B Multi-Agent Systems and Negotiation
Game theory provides foundational principles for automated negotiation. Nash bargaining [13] defines fair outcomes in bilateral settings. Rubinstein’s alternating offers model [16] describes sequential bargaining dynamics. These theoretical frameworks inform the protocol design presented in this work.
Recent multi-agent systems demonstrate potential for agent cooperation. MetaGPT [7] assigns distinct roles to collaborating agents, while CAMEL [10] facilitates communication through role-playing mechanisms. However, these systems primarily address collaborative task completion rather than bilateral negotiation involving competing interests. Privacy considerations are minimally addressed in their designs.
Zero-knowledge proofs facilitate private computation without disclosing underlying data [3]. zk-SNARKs [5] enable efficient proof generation suitable for practical deployment. Recent research applies these techniques to machine learning [9]. The present work integrates them into negotiation protocols. This allows agents to demonstrate constraint satisfaction without revealing the constraints themselves.
@@ II-C Explainability and Trust
Research in explainable AI demonstrates that audit trails improve user trust [14]. LIME and SHAP [11] provide post-hoc explanations for model predictions. Financial AI systems must comply with regulations that require decision transparency [4].
@@
Table I summarizes the comparison. The proposed framework combines autonomous reasoning, privacy-preserving negotiation, and cryptographic explainability within a device-native architecture.
System Capabilities Auto. Priv. Neg. Expl. Attest. Device State AutoGPT [15] ✓ – – – – – – LangChain [1] ✓ – – – – – MetaGPT [7] ✓ – – – – – CAMEL [10] ✓ – – – – – – Proposed Model ✓ ✓ ✓ ✓ ✓ ✓ ✓
@@ III-C Component 2: Explainable Memory
The explainable memory system produces cryptographic audit trails for each decision. Each record contains timestamp, decision type, inputs, reasoning trace, and outcome. These records are linked using Merkle trees [12], and the root hash is periodically anchored to a public blockchain to enable independent verification.
@@ III-D Component 3: World Model Distillation
Knowledge distillation with soft targets is applied [6]:
@@ III-E2 Negotiation Execution
Zero-knowledge proofs are constructed using Groth16 zk-SNARKs [5], which support succinct proofs with fast verification. The underlying circuit encodes constraint satisfaction, ensuring that the offered price meets , without revealing or . Proof generation requires 80 ms on high-end devices.
@@ III-H Integration and Data Flow
The architecture achieves its design objectives through component integration. Privacy is preserved through on-device execution, explainability through cryptographic audit trails, efficiency through selective offloading and model distillation, and safety through simulation-critic validation.
@@ IV-A Experimental Setup
Datasets. 2 public datasets are used for evaluation. The Medical Cost Personal dataset [2] contains 1,338 records with premium values ranging from $1,121 to $63,770, including features such as age, BMI, smoking status, and region. The Supply Chain Shipment Pricing dataset [18] comprises 10,324 B2B transactions with unit prices and features including product category, shipment mode, and freight costs.
@@ IV-B Component Performance
State Transfer. Table II presents selective state transfer results across device tiers. All configurations achieve over 70% compression while maintaining 85%+ task continuity, enabling seamless cross-device negotiation.
@@
Explainable Memory. Table III presents the memory system overhead and its impact on user trust. The system maintains low overhead across all operations.
@@
World Model Distillation. Table IV compares teacher and student models. Model size reduction enables device deployment. The 6% accuracy loss is acceptable given 5.5× latency improvement on high-end devices.
@@
Multi-Agent Negotiation. Table V presents the performance of the negotiation protocol. Success rates range from 85-90% across complexity levels with fairness scores of 0.83-0.88.
@@ IV-C End-to-End Performance
Insurance Negotiation. Table VI summarizes results for high-end devices across three complexity levels. The system achieves strong success rates while maintaining minimal privacy leakage compared to cloud approaches.
@@
B2B Negotiation. Table VII presents B2B procurement results. B2B scenarios demonstrate marginally better performance than insurance, with lower latency and comparable privacy preservation.
@@
Figure 3 shows success rate decline with complexity. Insurance exhibits steeper degradation (5 points) than B2B (4 points), reflecting its more contentious nature. Both domains maintain 83% success.
Figure 4 shows latency breakdown by component. The negotiation protocol dominates total time, followed by planning, demonstrating that multi-round bargaining is the primary latency driver.
@@ IV-D Baseline Comparison
The proposed system balances all metrics effectively. It achieves 2.4× speedup over Cloud-Only while maintaining strong privacy (94% reduction in data exposure) and highest success rates. Device-Only offers zero leakage but cannot handle complex reasoning, while Naive Edge shows intermediate performance across dimensions.
@@ IV-E Ablation Study
State Transfer Removal. Latency increases 108% as full state must be transmitted. Success drops 3% due to timeout failures.
World Model Removal. Cloud-based calls replace on-device planning, resulting in a 119% increase in latency and a 5% decrease in success rate. Privacy leakage also increases because planning data is transmitted to the cloud.
Negotiation Protocol Removal. Success rate decreases by 15% as agents cannot verify constraint satisfaction or engage in multi-round convergence. Latency improves by 17% because ZK proof generation and verification are eliminated. But this represents a loss of essential privacy-preserving functionality rather than a viable optimization.
@@
Figure 6 illustrates each component’s contribution to overall system performance. All components contribute meaningfully. The negotiation protocol and safety critic have the largest impact on success rates.
@@ V Discussion
Key Insights. Device-native agents demonstrate effective negotiation performance, matching or exceeding cloud-based alternatives while preserving privacy. Explainability significantly increases user trust with minimal overhead. Prior research confirms that transparency is crucial for AI systems in finance [19].
The proposed architecture is applicable beyond insurance and B2B domains. Core technical requirements such as bilateral bargaining [13, 16], privacy-sensitive constraints [3], and audit trails [4] are also present in real estate negotiations, healthcare cost discussions, and legal settlements. The cryptographic protocol and on-device execution model are domain-agnostic, requiring only the definition of domain-specific constraints and the training of the world model [6].
@@ VI Limitations
Experimental Scope. The evaluation was conducted in controlled scenarios [2, 18]. Real-world deployment studies are necessary to achieve more robust validation.
System Constraints. The system currently supports only bilateral negotiations [13, 16]. Multi-party scenarios and adversarial counterparties require additional safeguards. Device limitations restrict the reasoning depth of the 500M-parameter model [6] compared to cloud-scale models.
Behavioral Privacy Leakage. While the system achieves a 94% reduction in cryptographic data exposure, zero-knowledge proofs [3] do not prevent behavioral inference attacks [8]. Adversaries observing negotiation dynamics may infer constraint proximity. Current metrics quantify only cryptographic leakage (Sec. IV). Mitigation strategies, such as differential privacy for timing and randomized concession schedules, are identified as directions for future work.
Offline Collusion. The system cannot prevent offline collusion between parties who coordinate externally and use agents to generate audit trails. This is a social rather than technical limitation. However, cryptographic audit trails [12] can facilitate post-hoc fraud detection [17] to support regulatory compliance [4].